OpenStudy (anonymous):

What is the statement P(k + 1) for P(n): 2 + 4 + 8 + ... + 2n = 2(n + 1) - 2 I will give a medal!

OpenStudy (shubhamsrg):

Oh really?You;re gonna give medal? You;re so kind! :)

OpenStudy (anonymous):

P(n): 2 + 4 + 8 + ... + 2n = 2(n + 1) - 2 P(k): 2 + 4 + 8 + ... + 2k = 2(k + 1) - 2 P(1): 2 + 4 + 8 + ... = 2n = 2(1 + 1) - 2

OpenStudy (anonymous):

I just need the statement for P(k + 1).

OpenStudy (ash2326):

P(k): 2 + 4 + 8 + ... + 2k = 2(k + 1) - 2 replace k by k+1 P(k+1):2 +4+8+2(k+1)=2(k+1+1)-2

OpenStudy (aravindg):

replace n with k+1

OpenStudy (anonymous):

So P(k + 1): 2 + 4 + 8 + 2(k + 1) = 2(k + 1 + 1) - 2???

OpenStudy (ash2326):

yes, don't miss the .......

OpenStudy (anonymous):

Where would that go? P(k + 1): 2 + 4 + 8 + ... + 2(k + 1) = 2(k + 1 + 1) - 2 Is that right?

OpenStudy (aravindg):

yep :)

OpenStudy (anonymous):

Yayy! Thank you both so much! :) Have a nice day!

OpenStudy (ash2326):

:D

OpenStudy (aravindg):

you too :)